OpenStudy (anonymous):

26+6b ≥ 2(3b+4) i've been out of school for a year and i'm extremely rusty i apologize

OpenStudy (anonymous):

Are you trying to find b?

OpenStudy (anonymous):

it just says to solve it and graph it. if i could solve it, i'll have no trouble graphing it. that part i remember

OpenStudy (anonymous):

Ok so just think the inequality symbol as a = sign so then just solve for b.

OpenStudy (anonymous):

i got b=22..... i don't think that's right

OpenStudy (jdoe0001):

if you expand the right-hand side, what do you get?

OpenStudy (anonymous):

what

OpenStudy (jdoe0001):

if you were to expand say 2(3b+4), what would that expand to?

OpenStudy (anonymous):

oh, 6b+12

OpenStudy (jdoe0001):

hmmm there's no solution, as far as I can tell..... because anyhow 2 * 4 = 8 \(\ne 12\)

OpenStudy (jdoe0001):

\(\bf 26+6b\ge 2(3b+4) \implies 26\cancel{+6b}\ge \cancel{6b}+8\implies 26 \ge 8\)

OpenStudy (texaschic101):

26 + 6b >= 2(3b + 4) 26 + 6b >= 6b + 8 6b - 6b >= 8 - 26 0 >= - 18 (incorrect) I agree.....no solution

OpenStudy (jdoe0001):

hmmm.... actually 0 is greater than -18 I was wondering on that. because 26 is indeed greater than 8, I mean, is a valid inequality

OpenStudy (jdoe0001):

on the negative side, the closer to 0, the greater than number

OpenStudy (anonymous):

perfect, that's one of the choices. thank you guys so much

OpenStudy (texaschic101):

oops,,,my bad...I should have caught that 0 is greater then -18. So what is the answer ? 0b >= -18

OpenStudy (jdoe0001):

well. is valid... now.... how to graph it.... hmmm maybe just 2 dots =) one at 26 and one at 8 hehhe

OpenStudy (anonymous):

oh, well the answer was all real numbers i failed

OpenStudy (texaschic101):

well...yes because any number times 0 will equal 0, and that would be less then -18.

OpenStudy (texaschic101):

oops...greater

OpenStudy (anonymous):

I'M SO BAD AT THIS WOW thank you guys

OpenStudy (texaschic101):

you will get better...it just takes practice. And believe me when I tell you if I had a dollar for every mistake I made, I would be rich.